# PLANesT-3D: A new annotated dataset for segmentation of 3D plant point clouds

### v1

@@ 2.2 Reconstruction and pre-processing of 3D point clouds
2.3 The PLANesT-3D dataset
The PLANesT-3D dataset consists of 34 plant point clouds, where the plant points are isolated from the background as described in the previous section. Accompanying the locations and color information of the points in each model, three scalar fields are provided: The confidence map as delivered by Metashape Professional, semantic labels, and instance labels. Examples of point clouds from the PLANesT-3D dataset and their corresponding confidence maps and labels are depicted in Fig. 3.
Both semantic labels and instance labels are obtained through manually labeling each point using CloudCompare point cloud processing software [44]. Semantic labels for PLANesT-3D correspond to "leaf" and "stem" classes. Points on the leaf blades were labeled as "leaf", main stem, branches, and petioles were included in the "stem" class. Instance labels are identity numbers representing individual leaflets within a plant model. All points labeled as "stem" were assigned a single instance label.
@@ 2.3 The PLANesT-3D dataset
Table 3 gives five properties of each of the 34 plant point clouds in the dataset: 1) Plant height, 2) Total number of points, 3) Percentage of "leaf" points, 4) Percentage of "stem" points, and 5) Number of individual leaflets. Plant height is in the range of 25 to 105cm. The number of points exceed 1 million for most plants. Number of individual leaflets varies between 25 to 105, indicating the variation of complexity in plant structure.
@@ 2.4 Methods for semantic segmentation
Harandi et al. [22] provided a comprehensive review of typical steps involved in the processing and analysis of 3D representations of plants, including point clouds. Models in PLANesT-3D can be used to evaluate a diverse set of tools developed for plant analysis and phenotyping. In this study, we focus on the application of segmentation of 3D point clouds of plants into their semantic units.
@@
We developed a novel semantic segmentation method that is a combination of a superpoint extraction scheme [38] and an adaptive 3D object classification network [45]. We abbreviate this method as SP-LSCnet, where SP stands for superpoint, while LSCnet is a leaf-stem classification network. In order to provide reference performance results for further research, we also tested PointNet++ [46] and RoseSegNet [40] on the new PLANesT-3D dataset.
@@ 2.4.1 SP-LSCnet
Superpoint extraction: For accomplishing this task, the superpoint extraction scheme proposed in [38] is selected. The plant point cloud is first downsampled with voxel grid average filtering. t-SNE, as introduced in [47], is used to embed the downsampled 3D point cloud into 2D space. Superpoint extraction is then performed exclusively in 2D domain. t-SNE inherently provides clusters that are formed based on the implicit manifolds to which the 3D points belong.
@@
The first step of the superpoint extraction scheme corresponds to the identification of these clusters by Euclidean clustering in 2D. Line-like structures are detected through examining the local geometry of the points. An iterative procedure that involves spectral clustering and solidity computation is employed to segment non-convex clusters into convex regions. Once all 2D points are assigned to their corresponding superpoints, the assignments are carried to their 3D counterparts, and propagated to the high-resolution point cloud through nearest neighbor interpolation.
In this work, for the PLANesT-3D dataset, the grid size for voxel grid average filtering is selected as 0.12cm, the $Perplexity$ for t-sNE embedding is set to 30. The distance threshold used in Euclidean clustering in 2D space is chosen as 1. The other parameters were kept in their default settings as given in [38].
Figure 5: The integration of CSM and RUM to PointNet++ framework [45]: a) The classification network. $r^{L}$ for $L=1$ (first layer) is set to be 0.2. $r^{L}$ for $L=2$ (second layer) is 0.4. b) Adaptive local region inference with CSM and RUM.
@@
Leaf-stem classification network: Each superpoint generated by the previous step is considered to be a separate point cloud. A point-based deep neural network, introduced by [45], is utilized to classify the superpoints into leaf and stem classes (Fig. 4). The network is formed through integrating two modules to the PointNet++ architecture for object classification (Fig. 5). The modules are based on attention mechanisms to analyze point interactions within and between point neighborhoods, and are named as Center Shift Module (CSM) and Radius Update Module (RUM) [45]. Their task is to update adaptively the centers and radii of the spherical regions on which PointNet++ encodes local features. This operation is referred to as adaptive local region inference [45].
The classification network architecture of PointNet++ [46], shown in Fig. 5(a), includes three set abstraction (SA) layers. The input point cloud with $N=1024$ points is reduced to $N_{1}=512$, and $N_{2}=128$ representative points during SA layer-1 and SA layer-2, respectively. At each SA layer $L$, the representative points $c_{j}^{L}$ are selected using farthest point sampling (FPS) algorithm. $N_{L}$ spherical local regions with radius $r^{L}$ is defined around the center points $c_{j}^{L}$. Within each region, $K$ points are randomly selected and the features of $K$ points are mapped to higher dimensions with MLPs. The abstracted features of each local region are calculated by taking the maximum among the feature channels of $K$ grouped points.
The adaptive local inference modules, CSM and RUM can be integrated to either or both of the first two layers as shown in Fig. 5(b). If CSM is "ON" at a particular layer, for each representative point $c_{j}^{L}$, CSM calculates the shift amount $\Delta c_{j}^{L}$, through attention-based subnetworks. The representative point is updated as:
$\hat{c}_{j}^{L}=c_{j}^{L}+\Delta c_{j}^{L}\;.$ (4)
$\hat{r}_{j}^{L}=r^{L}+\Delta r_{j}^{L}\;.$ (5)
Various alternatives of CSM and RUM were suggested by [45]. As a result of experimental trials, the CSM-II (sub) and RUM-II (cum) versions were selected for the PLANesT-3D dataset. Please, refer to [45] for detailed descriptions of of the modules CSM-II (sub) and RUM-II (cum).
@@ 2.4.2 PointNet++
We adopted the PointNet++ architecture [46] for semantic segmentation of large point clouds. The data preparation for PointNet++ involves two operations: Subsampling and partitioning the big point set into subsets enclosed by blocks.
The subsampling is performed using voxel-grid filtering, where the edge length of the voxel cube is set as $1mm$. Each voxel is represented with a single point which is the average of the original points occupying the voxel. The sampled point is assigned the label of the closest point in the original cloud. This subsampling operation produces a point cloud with reduced size and homogeneous point density.
3D point-based deep neural networks requires input clouds with a fixed number of points. One way is to subsample the full plant point cloud to have the required number of points. However, such strategy will result in significant loss of geometric information, especially for small structures. Instead, we selected to partition the plant point cloud into subsets of points, each of which is then processed by the network individually. We followed a similar procedure of partitioning the point clouds into blocks as in [34, 40].
The horizontal region, i.e. the extent of the $XY$ plane, encompassing the point cloud is partitioned into squares of edge length of $10cm$. All points with $x$ and $y$ coordinates falling into a square region is considered belonging to the point subset within a block. The points in the block are subsampled to obtain a cloud of fixed number of points, as $N=8192$. The partitioning operation is conducted twice for each plant point cloud with offset values 00 and $5cm$.
We used the default PointNet++ architecture for semantic segmentation as given in [46]. PointNet++ segmentation network is composed of five grouping/abstraction layers. The radius parameter, which defines the size of local regions for grouping and abstraction, is set for each of the first four layers as $0.5cm$, $1cm$, $2cm$, and $4cm$, respectively. During training, batch size was selected as 16. The learning rate was updated as 0.005. Other hyperparameters of the PointNet++ were kept at default values.
@@ 2.4.3 RoseSegNet
RoseSegNet was developed for the specific application of semantic segmentation of rosebush models. It involves attention-based modules that encode point interactions within and between local regions. The details of the RoseSegNet architecture can be found in [40].
As with the case of PointNet++, we partition the plant point cloud into blocks and the point subset in each block is separately processed by RoseSegNet. The block partitioning process is the same as with PointNet++.
All the hyperparameters of the RoseSegNet are kept at their default values as given in [40], where the parameter search was performed on the ROSE-x dataset [26].
@@ 3 Results
In this section, we provide semantic segmentation results on the PLANesT-3D dataset obtained by the three methods described in Section 2.4. We trained a separate network for each of the three plant species. The hyperparameters of the networks were kept fixed for all the species. 70% of the point clouds were selected for training the networks. The rest is reserved for test. The split for training and test sets for the three species are given in Table 4. The IDs of the plants in the test sets are also provided.
@@
Before being processed by the networks, the point clouds were filtered such that points with confidence values less than 6 were discarded. For PointNet++ and RoseSegNet, the point cloud of a plant is partitioned into blocks and each block is separately processes as described in Section 2.4.2. For SP-LSCnet, the point clouds are partitioned into superpoints, and the superpoints are treated as separate objects for the object classification network. Table 5 gives the number of point subsets used for training and test for each plant species.
@@
We report the semantic segmentation results of the three methods in Table 6. Precision, Recall, and Intersection over Union (IoU) measures specific to "stem" and "leaf" classes are given in the table, as well as overall accuracy (Acc) and mean IoU (MIoU). In terms of overall accuracy and mean IoU, RoseSegNet yielded the best performance for all three species. For MIoU, RoseSegNet surpassed the performance of PointNet++ by close to 1%. SP-LSCnet is in between, performing slightly better than PointNet++ for the Pepper and Ribes sets, while yielding similar peformance to that of PointNet++ for the Rose set.
@@
The results on rose plant models are lower compared to those obtained on pepper and ribes plants. The recall of the "stem" class is more challenging for the rose plant models, since the petioles in between the leaflets are hard to discern. The misclassifications of petiole points of roses as belonging to leaf blades are observable in Fig. 6 and Fig. 7 where segmentation results are visually depicted for sample plants.
Figure 7: Examples to errors
@@
Other examples to erroneous cases are given in Fig. 7. The errors produced by PointNet++ and RoseSegNet appear as isolated points scattered in particular regions. Although context information heavily influences the features of single points in PointNet++ and RoseSegNet, the classification is performed individually for each point. Misclassifications occur particularly at cluttered regions where leaves and stems are close to each other.
@@
The treatment of each superpoint as a single entity by SP-LSCnet produces smooth results; however, it suffers from two disadvantages: First, the superpoints might not be always homogeneous; some leakage from other classes is inevitable due to the unsupervised nature of superpoint extraction. Second, the correct classification of the superpoints becomes particularly critical. As can be seen in Fig. 7, portions of the stem were classified as leaf points if the single superpoint to which they belong has the geometrical characteristics of leaves as modeled by the network. Likewise, small leaves, where the point resolution is not sufficient for proper sampling of the leaf surface, are classified as stem. Low point resolution due to occlusion (see bottom-left example in Fig. 7) can also cause isolated regions separated into superpoints which are not large enough to possess distinguishing shapes.

### v2

@@ 2.3 The PLANesT-3D dataset
The point clouds in the PLANesT-3D data possess color information as reconstructed by Metashape Professional. One important fact about the acquisition of the 2D images is that the illumination was not controlled. While all the camera images for an individual plant were acquired in one single session, image data from different plants were collected at different days and times of days. The illumination in the environment was not controlled. That is apparent from Fig. 4a, where apart from the natural variations in leaf texture, part of the color diversity comes from illumination changes, such as highlights and shadows due to the direction and amount of light.
As a result of these observations, we can conclude that the point clouds in PLANesT-3D dataset include leaf and stem instances of both high and low point density and quality with varying shape, texture and color characteristics. These characteristics make the PLANesT-3D a challenging dataset for modeling various variations in realistic data acquisition scenarios.
Both semantic labels and instance labels of the 34 plant point clouds were obtained through manually labeling each point using CloudCompare point cloud processing software [cloudcompare]. Semantic labels for PLANesT-3D correspond to "leaf" and "stem" classes. Points on the leaf blades were labeled as "leaf"; main stem, branches, and petioles were included in the "stem" class. Instance labels are identity numbers representing individual leaflets within a plant model. A single instance label is assigned to all points semantically labeled as "stem".
Table 3 gives five properties of each of the 34 plant point clouds in the dataset: 1) Plant height, 2) Total number of points, 3) Percentage of "leaf" points, 4) Percentage of "stem" points, and 5) Number of individual leaflets. Plant height is in the range of 25 to 105cm. The number of points exceed 1 million for most plants. Number of individual leaflets varies between 21 to 105, indicating the variation of complexity in plant structure.
Accompanying the locations and color information of the points in each model, three scalar fields are provided: The confidence map as delivered by Metashape Professional, semantic labels, and instance labels. Examples of point clouds from the PLANesT-3D dataset and their corresponding confidence maps and labels are depicted in Fig. 6.
@@
Figure 7: Percentage of remaining points as the confidence threshold varies.
Figure 8: The effect of confidence threshold.
Figure 9: Examples of raw point clouds with instance labels of the target plant.
As stated before, the confidence value for a point corresponds to the number of images where the point is observable, and that contribute to the 3D reconstruction of the point. It can be viewed as a measure of confidence on the accuracy of the estimated 3D location. The point cloud can be filtered by discarding points that have confidence values less than a certain threshold. Fig. 7 gives the box plots of the percentage of the number of points in a 3D model with confidence values above and equal to the confidence thresholds varying between 1 and 21. A separate box plot is given for each species. The bottom and top of each box are the 25th and 75th percentiles among the point clouds in the species, the whiskers indicate the range and the red line in the middle of each box gives the median value. We can observe a significant drop in the number of points as the threshold for the number of cameras observing the point increases. The drop in percentage is lower and less varied for the rose plants since the rose plants are smaller in size making it easier to cover the plant surface with more camera viewpoints observing smaller areas. More importantly, the rose plants suffer from self-occlusion to a lower degree compared to Pepper and Ribes.
Fig. 7 demonstrates the effect of filtering point clouds with varying confidence thresholds. As the number of camera viewpoints required to reconstruct a point increases, the highly visible parts remain relatively of high point density. The quality of broad leaves which are oriented perpendicularly to camera view orientations is not degraded. Rather, reconstruction noise on the surface of such leaves is reduced. However, around dense foliage, such as the top of the pepper and ribes plants, small leaves, and leaves oriented in parallel to the camera view orientations loose significant amount of points. Also, thin branches and petioles and branches and leaves hidden by heavy self-occlusion begin to disappear. The confidence values are provided as a scalar field of the models in the PLANesT-3D dataset for further investigation of the trade-off between reduced reconstruction noise and loss of self-occluded and fine-detail data.
In addition to the labeled 3D point clouds of individual plants, which were isolated from the background, we also provide the full raw point cloud as reconstructed by Metashape Professional. The instance labels were mapped onto the plant points in the scene. Examples of such point clouds are given in Fig. 9. These point clouds will be instrumental to test algorithms that automatically separate target plants from the environment. The 2D color images for all the 34 plants together with their estimated camera poses and parameters are also open to the public to provide input data for recent 3D reconstruction techniques [^3] The data is available at https://github.com/visionlab-ogu/PLANesT-3D/tree/main/data.
@@ 2.4 Methods for semantic segmentation
harandi2023make provided a comprehensive review of typical steps involved in the processing and analysis of 3D representations of plants, including point clouds. Models in PLANesT-3D can be used to evaluate a diverse set of tools developed for plant analysis and phenotyping. In this study, we focus on the application of segmentation of 3D point clouds of plants into their semantic units.
@@
We developed a novel semantic segmentation method that is a combination of a superpoint extraction scheme [dutagaci2023using] and an adaptive 3D object classification network [turgut2023local]. We abbreviate this method as SP-LSCnet, where SP stands for superpoint, while LSCnet is a leaf-stem classification network. In order to provide reference performance results for further research, we also tested PointNet++ [qi2017pointnet2] and RoseSegNet [Turgut2022rosesegnet] on the new PLANesT-3D dataset.
@@ 2.4.1 SP-LSCnet
Superpoint extraction: For accomplishing this task, the superpoint extraction scheme proposed in [dutagaci2023using] is selected. The plant point cloud is first downsampled with voxel grid average filtering. t-SNE, as introduced in [van2008visualizing], is used to embed the downsampled 3D point cloud into 2D space. Superpoint extraction is then performed exclusively in 2D domain. t-SNE inherently provides clusters that are formed based on the implicit manifolds to which the 3D points belong. The assumption that the data points lie on smooth low-dimensional manifolds accords with the fact that, for many plant species, plant points lie on one or two-dimensional manifolds. t-SNE makes it possible to unwrap the structure of the 3D plant model onto a single 2D view.
@@
The first step of the superpoint extraction scheme corresponds to 2D Euclidean clustering to separate the clusters formed by t-SNE. Line-like structures are detected through examining the local geometry of the points. An iterative procedure that involves spectral clustering and solidity computation is employed to segment non-convex clusters into convex regions. Once all 2D points are assigned to their corresponding superpoints, the assignments are carried to their 3D counterparts, and propagated to the high-resolution point cloud through nearest neighbor interpolation.
In this work, for the PLANesT-3D dataset, the grid size for voxel grid average filtering is selected as 0.12cm prior to t-SNE embedding. Grid size controls the computational load of t-SNE execution and the level of detail that is desired to be transferred to the 2D domain as the preserved local structure. For the plant models in the dataset, we aimed for a resolution close to 1mm while reducing the number of points to be processed by t-SNE. The $Perplexity$ value determines the effective neighborhood in terms of number of data points for t-SNE to calculate similarity. We targeted to have 3 to 4 cm of range for to measure the similarity between two points during the operation of t-SNE. Therefore, given that the grid size is set to 0.12cm, we chose the $Perplexity$ value as 30. The distance threshold used in Euclidean clustering in 2D space is chosen as 1 through the observation that minimum separation between clusters in 2D space is around one unit. The other parameters were kept in their default settings as given in [dutagaci2023using].
@@
Leaf-stem classification network: Each superpoint generated by the previous step is considered to be a separate point cloud. A point-based deep neural network, introduced by [turgut2023local], is utilized to classify the superpoints into leaf and stem classes (Fig. 10). The network is formed through integrating two modules to the PointNet++ architecture for object classification. The modules are based on attention mechanisms to analyze point interactions within and between point neighborhoods, and are named as Center Shift Module (CSM) and Radius Update Module (RUM) [turgut2023local]. Their task is to update adaptively the centers and radii of the spherical regions on which PointNet++ encodes local features. This operation is referred to as adaptive local region inference [turgut2023local]. The details of the network are provided in Appendix A.
@@ 2.4.2 PointNet++
We adopted the PointNet++ architecture [qi2017pointnet2] for semantic segmentation of large point clouds. The data preparation for PointNet++ involves two operations: Subsampling and partitioning the big point set into subsets enclosed by blocks.
The subsampling is performed using voxel-grid filtering, where the edge length of the voxel cube is set as $0.1cm$. Each voxel is represented with a single point which is the average of the original points occupying the voxel. The sampled point is assigned the label of the closest point in the original cloud. This subsampling operation produces a point cloud with reduced size and homogeneous point density.
3D point-based deep neural networks require input clouds with a fixed number of points. One way is to subsample the full plant point cloud to have the required number of points. However, such strategy will result in significant loss of geometric information, especially for small structures. Instead, we selected to partition the plant point cloud into subsets of points, each of which is then processed by the network individually. We followed a similar procedure of partitioning the point clouds into blocks as in [Turgut2022segmentation, Turgut2022rosesegnet].
The horizontal region, i.e. the extent of the $XY$ plane, encompassing the point cloud is partitioned into squares of edge length of $10cm$. All points with $x$ and $y$ coordinates falling into a square region is considered belonging to the point subset within a block. The points in the block are subsampled to obtain a cloud of fixed number of points, as $N=8192$. The partitioning operation is conducted twice for each plant point cloud with offset values 0 and $5cm$.
We used the default PointNet++ architecture for semantic segmentation as given in [qi2017pointnet2]. PointNet++ segmentation network is composed of five grouping/abstraction layers. The radius parameter, which defines the size of local regions for grouping and abstraction, is set for each of the first four layers as $0.5cm$, $1cm$, $2cm$, and $4cm$, respectively. During training, batch size was selected as 16. The learning rate was updated as 0.005. Other hyperparameters of the PointNet++ were kept at default values.
@@ 2.4.3 RoseSegNet
RoseSegNet was developed for the specific application of semantic segmentation of rosebush models. It involves attention-based modules that encode point interactions within and between local regions. The details of the RoseSegNet architecture can be found in [Turgut2022rosesegnet].
As with the case of PointNet++, we partitioned the plant point cloud into blocks and the point subset in each block was separately processed by RoseSegNet. The block partitioning process is the same as with PointNet++.
All the hyperparameters of the RoseSegNet are kept at their default values as given in [Turgut2022rosesegnet], where the parameter search was performed on the ROSE-X dataset [Dutagaci2020rose].
@@ 3 Results
In this section, we provide semantic segmentation results on the PLANesT-3D dataset obtained by the three methods described in Section 2.4. We trained a separate network for each of the three plant species. The hyperparameters of the networks were kept fixed for all the species. 70% of the point clouds were selected for training the networks. The rest is reserved for test. The splits for training and test sets for the three species are given in Table 4. The IDs of the plants in the test sets are also provided.
@@
Before being processed by the networks, the point clouds were filtered such that points with confidence values less than 6 were discarded. For PointNet++ and RoseSegNet, the point cloud of a plant is partitioned into vertical blocks of size $10cm\times 10cm$ of base area, with offset values of 0 and $5cm$. The points in each block are subsampled to obtain $N=8192$ points per block. Each block is processed by the PointNet++ and RoseSegNet independently as a separate point cloud, as described in Section 2.4.2. The number of blocks extracted from the training and test plant models are given in Table 5. The training blocks correspond to the blocks extracted from the plant point clouds reserved for training the networks.
For SP-LSCnet, the point clouds are partitioned into superpoints, and the superpoints are treated as separate objects for the object classification network. Table 5 gives the number of superpoints extracted from the training and test plant models for each plant species.
@@
Figure 11: Training curves for the classification network LSCnet for the three species. Accuracy corresponds to the classification accuracy of superpoints in training and evaluation (test) sets.
The classification network LSCnet takes superpoints as individual point clouds as input and trains the network to classify the superpoints into "leaf" or "stem" objects. A separate model was trained for each of the three species. Fig. 11 gives the training curves for the classification network LSCnet for the three species. Accuracy corresponds to the classification accuracy of the superpoints in training and evaluation (test) sets. Maximum number of epochs was set to 250. The accuracy on the training set continued to increase steadily. The accuracy on the test set kept increasing, although slightly and with oscillations, until 150th epoch for the pepper plants and until 250th epoch for the rose plants. For the ribes plants, the performance on the test set did not improve further after 20 epochs. However, the oscillations did not fall below the accuracy rate of 0.9 in the subsequent epochs. These results indicate that the LSCnet did not suffer from overfitting during the training process.
@@ 3.1 Quantitative results
We report the semantic segmentation results of the three methods in Table 6. Precision, Recall, and Intersection over Union (IoU) measures specific to "stem" and "leaf" classes are given in the table, as well as overall accuracy (Acc) and mean IoU (MIoU). In terms of overall accuracy and mean IoU, RoseSegNet yielded the best performance for all the three species. In terms of MIoU, RoseSegNet surpassed the performance of PointNet++ by close to 1%. SP-LSCnet is in between, performing slightly better than PointNet++ for the Pepper and Ribes sets, while yielding similar peformance to that of PointNet++ for the Rose set.
Figure 13: Examples of errors of the three methods on Pepper point clouds.
Figure 14: Examples of errors of the three methods on Rose point clouds.
Figure 15: Examples of errors of the three methods on Ribes point clouds.
@@
The segmentation performance figures on rose plant models are lower compared to those obtained on pepper and ribes plants. The recall of the "stem" class is more challenging for the rose plant models, since the petioles in between the leaflets are hard to discern. For other plants, the MIoU figures are still below 96 %, indicating that there is room for improvement. The following subsection, where a visual error analysis is given, provides more insight into the causes of errors.
3.2 Visual results
Fig. 12 gives visual semantic segmentation results by PointNet++, RoseSegNet, SP-LSCnet on sample point clouds. In Figs. 13, 14, and 15 examples to erroneous regions produced by the three methods are demonstrated for pepper, rose, and ribes plants, respectively. For the rose plants, all three methods produce erroneous results at thin petioles where the opposing leaflets are very close (Fig. 14). The misclassified points of PointNet++ and RoseSegNet are observed as speckles. With SP-LSCnet, the erroneous regions are smooth as they were misclassified together with the superpoint they belong to.
@@ 3.2 Visual results
The methods also produce errors at leaves and other stem parts of the plants; however the nature of the errors of SP-LSCnet is different from those of PointNet++ and RoseSegNet. With the laters, the errors appear as isolated points scattered in the regions (Figs. 13, 14, and 15). Although context information heavily influences the features of single points in PointNet++ and RoseSegNet, the classification is performed individually for each point. Misclassifications occur particularly at cluttered regions where leaves and stems are close to each other.
@@
The treatment of each superpoint as a single entity by SP-LSCnet produces smooth results; however, it suffers from two disadvantages: First, the superpoints might not be always homogeneous; some leakage from other classes is inevitable due to the unsupervised nature of our approach for superpoint extraction. Second, the correct classification of the superpoints becomes particularly critical. At the junctions, portions of stems may be separated from the adjacent branches and misclassified as leaves, as observed in Fig. 13c and Fig. 15c. In these cases, the stem points were classified as leaf points since the single superpoints to which they belong have the geometrical characteristics of leaves as modeled by the network. Likewise, small leaves, where the point resolution is not sufficient for proper sampling of the leaf surface, are classified as stem (Fig. 13c).
PointNet++ and RoseSegNet have a tendency to misclassify borders of leaves as stems as can be obserfed in Figs. 13, 14 and 15. Such errors are rare with SP-LSCnet since borders of a leaf are not partitioned into separate superpoints. However, when there is a hole in the leaf or the point density drops at a certain region, as are the two cases in the leaves shown in Fig. 13(c), parts of the leaves can be separated into different superpoints and misclassified as stems.
An important cause of error of the SP-LSCnet is low point density and missing structures due to self occlusion. Regions of low point density can cause isolated regions separated into superpoints which are not large enough to possess distinguishing shapes. Three examples of such errors are illustrated in Fig. 15(c).